OpenStudy (anonymous):

Use the quadratic formula to solve (x − 4)(x + 5) = 7.

OpenStudy (shane_b):

So where would you start?

OpenStudy (alexwee123):

(x − 4)(x + 5) = 7. x^2 +x-20=7 x^2 +x-27=0  can you solve it out now? @jorgereza9